TEX Users Group

Technical Council

Technical Council Members:

Kaja P. Christiansen (Chair)

kaja@daimi.aau.dk

Michael Ferguson

mike@inrs-telecom.uquebec.ca

Yannis Haralambous

Yannis.Haralambous@univ-lille1.fr

Sebastian Rahtz

s.rahtz@elsevier.co.uk

Forming a Technical Working Group:

Here is a list of the currently active Technical Working Groups (TWGs). If you wish to form a TWG for an area of particular interest you should contact any one of the Technical Council members with your suggestion. We encourage any new TWG for which there is an enthusiastic chair and an interested constituency. If you wish to participate in a particular TWG, please contact the chair. All participation in a TWG is at the discretion of the TWG chair.

Technical Working Groups:

There are several types of TWGs:

IRP Independent Research Projects SC Standards and Coordination SI Special Interest

Each TWG name is prefixed by one of these types.

WG-92-00 (IRP-TWG) Title: Independent Projects Mandate: To recognize and report to the TEX Board and

the TEX Community on important projects which are independent of TUG but are of concern to the entire TEX Community. Due to its independent nature, each such project will retain complete control over its membership, working methods and agenda. The chair of the Independent Projects Working Group shall report, with the cooperation of a project’s management, on that project’s activities to the TUG Board and thence to the TEX Community. The reporting function of the TWG Chair shall in no way interfere or preclude reporting or dissemination of the results of a project by the project management or by members of the project.

TWG Chair: Alan Hoenig

ajhjj@cunyvm.cuny.edu

TC Liaison: Michael Ferguson WG-92-01 (SC-TWG) Title: TEX Extended

Mathematic Font Encoding Mandate: To create font encoding standards for Mathe-

matical fonts used in TEX systems.

TWG Chair: Barbara Beeton

            bnb@math.ams.com

TC Liaison: Michael Ferguson

WG-92-03 (SC-TWG) Title: Multiple Language

Coordination Mandate: The primary purpose of this working group is

to obtain, for TEX systems, a consistent means for implementing, accessing, and describing, the fonts, ligature rules, hyphenation patterns and other special requirements for a given linguistic group. Each linguistic group shall have a Group Coordination Chair who shall be responsible for discussions and implementation of those features and macros required for their group. In addition, the working group shall strive to create uniform mechanisms for access and definitions that are common to each linguistic group.

TWG Chair: Yannis Haralambous

Yannis.Haralambous@univ-lille1.fr

Common Language Issues: Johannes Braams

IL-Braams@pttnrl.nl

Current Group Coordination Chairs:

Group Coordination Chair (Greek): Silvio Levy

levy@math.berkeley.edu

Group Coordination Chair (Scandinavian): Jan Michael Rynning

jmr@nada.kth.se

Group Coordination Chair (African): Jörg Knappen

knappen@vkpmzd.kph.uni-mainz.edu

WG-92-04 (SI-TWG) Title: TEX for the Disabled Mandate: The primary purpose of this working group

is as a forum for those people interested in using and/or enhancing TEX to serve the needs of the disabled.

TWG Chair: T.V. Raman

            raman@crl.dec.com

TC Liaison: Michael Ferguson WG-92-05 (SC-TWG) Title: TEX Archive

Guidelines Mandate: The purpose of this Technical Working Group

is to develop guidelines for the effective management and utilization of major TEX archives, and to initiate communication among the maintainers of the existing archives for the purpose of coordination and synchronization. Areas to be considered include, but are not limited to:

– what physical resources are needed

– what kind of management support is desireable – what personnel resources are necessary and how

should they be organized

– breadth and depth of coverage, general and specific subject archives

– strategies for organizing and managing the archive holdings

– strategies for archive synchronization

– dissemination strategies of archive contents – should there be “official” archives

TWG Chair: George Greenwade

            bed gdg@shsu.edu

TC Liaison: Michael Ferguson

WG-93-06 (SI-TWG) Title: Visual typesetting and

document preparation Mandate: To investigate, compare and evaluate software

methods that allow a user to produce and alter a document by interacting with a graphic rather than ASCII representation. Emphasis will be on interaction between the editing and typesetting modules, and particularly on the requirements they place on each other. Although a working system is beyond the scope of the group, practical matters will be discussed.

TWG Chair: Jonathan Fine

            J.Fine@pmms.cam.ac.uk

TC Liaison: Yannis Haralambous

WG-94-07 (SC-TWG) Title: TEX Directory

Structures Mandate: The primary purpose of this TWG is to iden-

tify a universal directory structure for macros, fonts and other related TEX software so that recommendations can be made to all suppliers of TEX software.

TWG Chair: Norm Walsh

            norm@ora.com

TC Liaison: Sebastian Rahtz WG-94-08 (SC-TWG) Title: DVI Driver Implementation and Standardisation Issues Mandate: The major objective shall be to study the is-

sues in the requirements of DVI drivers imposed by changing needs and technologies, and to make recommendations for implentation and standardisation of such drivers to enhance the uniformity of their use. Work will include, but not be limited to, the examination of the use, syntax, and semantics of \special{..} commands.

TWG Chair: Michael Sofka

                                                              sofkam@rpi.edu

TC Liaison: Sebastian Rahtz

WG-94-09 (SI-TWG) Title: TEX and SGML Mandate: The major objective would be to investigate

the requirements and difficulties in developing an interface technology for TEX and SGML.

TWG Chair: Ken Dreyhaupt

                                                              kend@springer-ny.com

TC Liaison: Michael Ferguson

WG-94-10 (SI-TWG) Title: TEX and Linguistics Mandate: The main goal is to study and discuss the re-

quirements for typesetting linguistics in TEX and as a means of identifying, examining, testing, and comparing macros, fonts, style files and other aids for typesetting linguistics.

TWG Chair: Christina Thiele

                                                              cthiele@ccs.carleton.ca

TC Liaison: Yannis Haralambous

Source: /tex-archive/usergrps/tug/twg-list.tex on CTAN. Rev. Feb. 98